Vector diagrams
While snorkeling in the ocean, you swim directly towards shore <towards south >at 2m/s.The current of the water pushes you directly sideways <towards West >at 3m/s. Determine the magnitude and direction of the resultant velocity
1
Expert's answer
2019-01-17T08:19:11-0500

The magnitude of the resultant velocity

v=√(v_1^2+v_2^2 )=√(2^2+3^2 )≈3.6 m/s.

The direction of the resultant velocity

θ=tan^(-1)⁡〖v_2/v_1 〗=tan^(-1)⁡〖3/2〗=56° from the direction of the shore.